A soda machine contains $3.00 in nickels and dimes. If the number of dimes is 5 more than twice the number of nickels, how many coins of each type are there?

I am completely at a loss for understanding how to solve word problems. Can you help, please? Thank you!

Answer by Alan3354(69443)

A soda machine contains $3.00 in nickels and dimes. If the number of dimes is 5 more than twice the number of nickels, how many coins of each type are there?
n = # of nickels
d = # of dimes
------------
Write down what's given. Don't whine about "word problems"
"Two plus two" is a word problem.
--------------
d = 2n + 5 - # of dimes = 2 times # of nickels plus 5
5*n + 10*d = 300 - n's are 5 cents each, and dimes are 10 cents each. See that?
-----------------
d = 2n + 5
5*n + 10*d = 300
Now it's not a word problem.